### v1

@@ Appendix A Potential Limitation and Societal Impact
A.1 Contribution Statement
This work is the result of a large and comprehensive project that aims to provide the community with an open-source tool and platform for joint understanding and generation of sounding video, which is the first of its kind. While most of the individual network components are known in literature, it takes significant trials and errors to find a feasible plan, and connect and align them to form a state-of-the-art and yet concise system. In particular, the proposed three-stage training pipeline has proven useful for gradually improving model performance. In addition, the proposed JavisInst-Omni dataset represents a very rich mixture of single- and multi-modality data, with the majority being single-turn and the rest multi-turn. We believe the architecture, training methods, and datasets offer a very competitive baseline and have significant value for future data-centric and model-centric studies.
A.2 Limitation and Future Work
@@
Architecture. JavisGPT adopts a straightforward encoder–LLM–decoder architecture. While effective, as briefly mentioned in Sec.˜6, this design presents two core inconsistencies: (1) Misaligned training objectives. The comprehension tasks rely on next-token prediction (NTP) loss [bai2025qwen25vl] for textual outputs, whereas the generation tasks use diffusion loss [liu2025javisdit] (with alignment loss) for audio-visual outputs. These distinct objectives lead to inconsistent optimization signals for the shared LLM backbone. (2) Asymmetric input–output modeling. For multimodal content (in this case, audio and video), the input relies on VAE-based continuous embeddings [bai2025qwen25vl], while the output is conditioned through a query-based interface [pan2025metaquery]. The connection between the two is indirect: the generation queries can “see” the embeddings from the comprehension stage, allowing understanding to enhance generation; however, the reverse is not true, as the generation queries are not part of the comprehension input. To achieve a truly unified MLLM for audio-video understanding and generation, it may be beneficial to draw inspiration from recent advances in unified modeling for image-based multimodal tasks [wang2024emu3, wu2024vila, tong2024metamorph] and adapt them to the audio-video domain. Appendix˜D further elaborates on possible architectural directions for unifying comprehension and generation, as well as improving spatiotemporal synchronization, which we hope can provide valuable insights for future research in the community.
@@ B.3 Evaluation Setup
In Sec.˜5, we comprehensively evaluate the multimodal understanding and generation performance of JavisGPT, and here we introduce the detailed evaluation setups.
@@ D.1 Design Choices of Generative MLLM
In conclusion, the LLM+DiT paradigm is the best option for current JAVG field, which has also been verified by our comprehensive experiments in Sec.˜5. However, the LLM+AR+Decoder architecture may show potentials to further bridge the unified task of comprehension and generation. The key insight is simple and straightforward: after unifying understanding and generation AV representations into the same space, the generation embeddings can be also “seen” by comprehension tasks. As discussed in Appendix˜A, this feature can mutually strengthen both comprehension and generation capabilities after scaling up model capacity and data quantity. We hope this can bring more inspiration in the community.
@@ D.2 Design Choices of LLM-DiT Combination
NExT-GPT. NExT-GPT [wu2024next] formulates generation as a classification task over a fixed vocabulary of video IDs (e.g., <VID1>, <VID2>, $\cdots$, <VIDn>), forcing all instruction prompts to regress to this small set of special tokens. This rigid formulation limits the expressiveness of conditional instructions and severely constrains the diversity and fidelity of generated audio-video outputs, which is consistent to NExT-GPT’s poor generation performance in Sec.˜5.1. Additionally, it restricts generalization to novel or fine-grained instructions outside the pre-defined token set.
@@ E.1 Details for Human Evaluation on Interleaved Conversation
In Sec.˜5.1, we present a comparative evaluation with UnifiedIO-2 [lu2024unifiedio2] and NExT-GPT [wu2024next] in interleaved comprehension and generation scenarios. Since no existing benchmark is available for this setting, we conduct a human evaluation, as detailed below:

### v2

@@ Appendix A Potential Limitation and Societal Impact
A.1 Limitation and Future Work
@@
Architecture. JavisGPT adopts a straightforward encoder–LLM–decoder architecture. While effective, as briefly mentioned in Sec.˜7, this design presents two core inconsistencies: (1) Misaligned training objectives. The comprehension tasks rely on next-token prediction (NTP) loss [bai2025qwen25vl] for textual outputs, whereas the generation tasks use diffusion loss [liu2025javisdit] (with alignment loss) for audio-visual outputs. These distinct objectives lead to inconsistent optimization signals for the shared LLM backbone. (2) Asymmetric input–output modeling. For multimodal content (in this case, audio and video), the input relies on VAE-based continuous embeddings [bai2025qwen25vl], while the output is conditioned through a query-based interface [pan2025metaquery]. The connection between the two is indirect: the generation queries can “see” the embeddings from the comprehension stage, allowing understanding to enhance generation; however, the reverse is not true, as the generation queries are not part of the comprehension input. To achieve a truly unified MLLM for audio-video understanding and generation, it may be beneficial to draw inspiration from recent advances in unified modeling for image-based multimodal tasks [wang2024emu3, wu2024vila, tong2024metamorph] and adapt them to the audio-video domain. Appendix˜D further elaborates on possible architectural directions for unifying comprehension and generation, as well as improving spatiotemporal synchronization, which we hope can provide valuable insights for future research in the community.
@@ B.3 Evaluation Setup
In Sec.˜6, we comprehensively evaluate the multimodal understanding and generation performance of JavisGPT, and here we introduce the detailed evaluation setups.
@@ D.1 Design Choices of Generative MLLM
In conclusion, the LLM+DiT paradigm is the best option for current JAVG field, which has also been verified by our comprehensive experiments in Sec.˜6. However, the LLM+AR+Decoder architecture may show potentials to further bridge the unified task of comprehension and generation. The key insight is simple and straightforward: after unifying understanding and generation AV representations into the same space, the generation embeddings can be also “seen” by comprehension tasks. As discussed in Appendix˜A, this feature can mutually strengthen both comprehension and generation capabilities after scaling up model capacity and data quantity. We hope this can bring more inspiration in the community.
@@ D.2 Design Choices of LLM-DiT Combination
NExT-GPT. NExT-GPT [wu2024next] formulates generation as a classification task over a fixed vocabulary of video IDs (e.g., <VID1>, <VID2>, $\cdots$, <VIDn>), forcing all instruction prompts to regress to this small set of special tokens. This rigid formulation limits the expressiveness of conditional instructions and severely constrains the diversity and fidelity of generated audio-video outputs, which is consistent to NExT-GPT’s poor generation performance in Sec.˜6.1. Additionally, it restricts generalization to novel or fine-grained instructions outside the pre-defined token set.
@@ E.1 Details for Human Evaluation on Interleaved Conversation
In Sec.˜6.1, we present a comparative evaluation with UnifiedIO-2 [lu2024unifiedio2] and NExT-GPT [wu2024next] in interleaved comprehension and generation scenarios. Since no existing benchmark is available for this setting, we conduct a human evaluation, as detailed below: